OpenStudy (anonymous):

Can someone work through it with me? What is the value of (7 – 3)!?

OpenStudy (anonymous):

7-3 7 6 5 4 7-3=4

OpenStudy (mathmale):

Hint: 7-3 = ? Hint #2: 4! = 4*3*2*1 = ?

OpenStudy (anonymous):

That's not exactly what I meant it's not 7-4

OpenStudy (anonymous):

24

OpenStudy (mathmale):

I'm afraid we'll need to start over, then. What was your question about?

OpenStudy (anonymous):

it's coutning and probabilty

OpenStudy (anonymous):

What is the value of (7 – 3)!? A. 4 B. 24 C. 840 D. 5036

OpenStudy (mathmale):

(7-3)! has a specific meaning in algebra. I asked you to calculate 7-3 first, and then use the result in evaluating ( ) ! 7-3=4, as you said; so, what is 4! ? Hint: 4! is read "four factorial." I have already shown you how to calculate it.

OpenStudy (mathmale):

By the way, I really like your introduction, "Can somebody work through this with me?"

OpenStudy (anonymous):

4*3*2*1= 24 And thank you. :)

OpenStudy (mathmale):

Nice work! Our conversation here was so much more satifying and proper than was the previous one. Thanks for the great attitude! Good luck.

OpenStudy (anonymous):

Thank you kind sir.